OpenStudy (ksaimouli):

find all values root2 tanx cosx - tanx=0 (0,2pie)

OpenStudy (amistre64):

use () to group that arguments under the root2

OpenStudy (ksaimouli):

ya i got the answer as tan x=0 and cosx 1/root2

OpenStudy (mr.math):

Is it \(\sqrt{\tan{x}\cos{x}-\tan{x}}=0\)?

OpenStudy (mr.math):

Or \(\sqrt{\tan{x}\cos{x}}-\tan{x}=0\)?

OpenStudy (ksaimouli):